# Trajectory Volatility for Out-of-Distribution Detection in Mathematical Reasoning

### v1

@@ 2.2 Early Stabilization: A Micro Empirical Perspective
We refer to this phenomenon as Early Stabilization under mathematical reasoning: For ID data, GLMs largely complete their reasoning in the mid-to-late stages, and simple adjustments are sufficient after that. However, for OOD data, GLMs can still not complete accurate reasoning at a later stage. They thus can only randomly switch to a specific output pattern, i.e., the scalar immediate number pattern. This enables an intuition that leverages the trajectory for OOD detection may be effective.
3 TV Score: Trajectory Volatility for OOD Detection
In this section, we propose our trajectory-based algorithm TV Score to detect OOD samples in mathematical reasoning scenarios. Formally, for each input sample sequence ${\boldsymbol{x}}=x_{1}x_{2}...x_{n}$, the hidden embedding of $i$-th layer corresponding to the $i$-th token $x_{i}$ is denoted as ${\boldsymbol{h}_{i}^{l}}\in\mathbb{R}^{d}$. We set $n_{l}$ as the sequence length in $l$-th layer, and ${\boldsymbol{y}_{l}}$ to be the average embedding in the $l$-th layer, namely
${\boldsymbol{y}_{l}}=\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}{\boldsymbol{h}_{i}^{l}}~{}~{}~{}~{}(1\leq i\leq n_{l},1\leq l\leq L).$ (4)
Next, we need to obtain the trajectory volatility $V_{l}$ in each layer, then computer the final TV Score as:
$\textrm{TV}:=\frac{1}{L-1}\sum_{l=1}^{L-1}V_{l}.$ (5)
Therefore, defining the trajectory volatility $V_{l}$ stands as a key concern. The volatility between each neighboring layer-pair can be intuitively defined by the difference between the two embedding vectors, so the L1-norm $\|\boldsymbol{y}_{l}-\boldsymbol{y}_{l-1}\|_{1}$ is the simplest way. However, anomalous changes in each dimension under vector volatility will greatly affect the measure of volatility, introducing potentially uncontrollable factors. To address the challenge, we want to find the most critical principal component that maps the vector in low dimensions.
We consider using Mahalanobis Distance (MaDis) [26] to complete this mapping. We follow the assumptions of previous work [3, 46] and fit all average embedding ${\boldsymbol{y}_{l}}$ of ID samples in each layer to a Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l},{\boldsymbol{\Sigma}}_{l})$. For a new sample, We measure the MaDis between its hidden embedding ${\boldsymbol{y}}_{l}$ of $i$-th layer and $\mathcal{G}_{l}$, the conversion function $f({\boldsymbol{y}_{l}}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is
$f(\boldsymbol{y}_{l})=\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)~{}~{}~{}~{}(1\leq l\leq L)$ (6)
The underlying assumption is that when the MaDis between embeddings and corresponding Gaussian distribution of ID data in one layer and that of the neighboring layer has a larger difference, the embedding shift between two layers is more volatile. In this case, $V_{l}$ is defined as:
$V_{l}=\left|f(\boldsymbol{y}_{l})-f(\boldsymbol{y}_{l-1})\right|.$ (7)
Differential Smoothing (DiSmo).
@@ 3 TV Score: Trajectory Volatility for OOD Detection
Outliers in the trajectory may significantly impact feature extraction [25, 32] then lead to errors. To mitigate this, we explore higher-order differential smoothing techniques to enhance trajectory smoothness. We first define the Gaussian distribution difference and embedding difference based on the additivity of the Gaussian distribution as
$\mathcal{N}({\boldsymbol{\mu}}_{l}^{(k)},{\boldsymbol{\Sigma}}_{l}^{(k)})=\Delta^{(k)}\mathcal{G}_{l}$ $=\Delta^{(k-1)}\mathcal{G}_{l+1}-\Delta^{(k-1)}\mathcal{G}_{l}=\Delta^{(k-2)}\mathcal{G}_{l+2}-2\cdot\Delta^{(k-2)}\mathcal{G}_{l+1}+\Delta^{(k-2)}\mathcal{G}_{l}$ (8)
$=\cdots=\mathcal{N}(\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\boldsymbol{\mu}}_{l+k},~{}\sum_{i=0}^{k}\mathrm{C}_{k}^{i}{\boldsymbol{\Sigma}}_{l+k}),$
${\boldsymbol{y}_{l}^{(k)}}={\boldsymbol{y}_{l+1}^{(k-1)}}-{\boldsymbol{y}_{l}^{(k-1)}}=\cdots=\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\boldsymbol{y}}_{l+k},$ (9)
where $\mathrm{C}_{k}^{i}=\frac{k!}{i!(k-i)!}$. Similarly, we computer $k$-th order embedding difference MaDis:
$f^{(k)}(\boldsymbol{y}_{l})=\left[{\boldsymbol{y}_{l}^{(k)}}-{\boldsymbol{\mu}_{l}^{(k)}}\right]^{\top}\left[\Sigma_{l}^{(k)}\right]^{-1}\left[{\boldsymbol{y}_{l}^{(k)}}-{\boldsymbol{\mu}_{l}^{(k)}}\right],$ (10)
and in this case, $V_{l}$ is defined as:
$V_{l}=\left|f^{(k)}(\boldsymbol{y}_{l})-f^{(k)}(\boldsymbol{y}_{l-1})\right|.$ (11)
The complete pseudo-code is presented in Appendix C. In experiments, we choose $k\leq 5$ because if $k$ is too large, the over-smoothing phenomenon may happen and effective features will be eliminated.
@@ 4.1 Setup
For the ID dataset, we use the MultiArith [47], which consists of Math Word Problems on arithmetic reasoning. For the OOD datasets, we intuitively introduce two types of detection scenarios following [46]: (i) Far-shift OOD setting, we select the MATH [13] dataset with five domains of algebra, geometry, counting and probability, number theory, and precalculus; (ii) Near-shift OOD setting, we select five independent datasets: GSM8K [7], SVAMP [42], AddSub [15], SingleEq [21], and SingleOp [21]. We consider the ID data negative(-) and the OOD data positive(+). Refer to Appendix D.1 for basic information and OOD features of these datasets.
@@
Given the limited data size of MultiArith, totaling only 600 samples and lacking a standard division, we allocate 360 samples for training and 240 for testing. However, with such a small test set, randomness in evaluation becomes a concern. To mitigate this, we conduct test sampling and set the sampling size as 1000. Specifically, we denote ID dataset as $\mathcal{D}_{\text{in}}$ and OOD dataset as $\mathcal{D}_{\text{out}}$. For each sampling, the collection is $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$ where $\widetilde{\mathcal{D}}_{\text{out}}\subset\mathcal{D}_{\text{out}}$ and $|\mathcal{D}_{\text{in}}|=|\widetilde{\mathcal{D}}_{\text{out}}|$, this guarantees positive and negative sample balance. We report both the mean and standard variance of the results to enhance the reliability of evaluations. Refer to Appendix D.2 for the ID dataset split.
@@
To measure the application value of our method used in cutting-edge GLMs, we use Llama2-7B [53] and GPT-2-XL (1.5B) [6] as our backbones for ID dataset training. Refer to Appendix D.3 for training details. However, there exists uncertainty about the data used in the pre-training phase, especially for Llama2 because its data is closed-source. Some research [57, 59] have confirmed the absence of data leakage in Llama2 for the MATH and GSM8K datasets, we still conduct pre-experiments to examine the rationality of the OOD data selection rigorously. Our criterion is the claim that a dataset can be categorized as OOD if it exceeds the capabilities of the base model, as proposed in prior studies [50, 31]. Results of the pre-experiments are shown in Appendix D.4, and they can confirm that these datasets can be considered as OOD data for the two GLMs.
@@
We compare our method with some training-free baselines [^1] We refer to the latest survey [23] to select baselines for the scarcity of OOD detection methods on GLMs. where OOD training data are unavailable: (1) Maximum Softmax Probability (MS-Prob) [14]; (2) Monte-Carlo Dropout (MC-Drop) [10]; (3) Sequence Perplexity (PPL) [3]; (4) Input Embedding (I-Emb) [46]; (5) Output Embedding (O-Emb) [46]. Additionally, we set the smoothing order $k$ ranges from 1 to 5 and report the highest among them when with smoothing. Refer to Appendix D.5 for baseline details.
@@
We establish two main evaluation scenarios: (i) The basic OOD detection scenario, we further divide it into offline and online discriminations; (ii) More challenging and fine-grained OOD quality estimation. Task-specific metrics are presented in the corresponding result sub-sections.
4.2 Scenario I: OOD Detection
Offline Discrimination.
We utilize TV scores for offline OOD discrimination. For each collection $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$, we report the AUROC [39] and FPR95 metrics. The former represents the area under the ROC curve and the latter represents the value of FPR at 95% TPR. Table 1 presents the results.
@@ 4.2 Scenario I: OOD Detection
In addition, we conduct significant tests (Details are shown in Table 8 - 11). We find that our methods almost pass all significance tests, while the embedding-based methods have the lowest pass rate among baselines, suggesting that their results are more susceptible to sampling error.
@@
MS-Prob [14] 78.66$\pm$1.38 81.44$\pm$3.56 60.14$\pm$1.54 88.91$\pm$2.41 70.54$\pm$1.42 78.29$\pm$2.02 67.12$\pm$1.20 76.27$\pm$2.66
MC-Drop [10] 68.63$\pm$2.21 87.04$\pm$4.88 52.33$\pm$2.21 91.92$\pm$1.89 66.18$\pm$1.87 84.69$\pm$1.65 63.54$\pm$1.72 78.08$\pm$2.50
PPL [3] 85.64$\pm$1.46 53.06$\pm$4.36 59.35$\pm$1.89 86.09$\pm$1.89 80.82$\pm$1.04 64.53$\pm$2.10 73.74$\pm$1.12 72.39$\pm$1.27
I-Emb [46] 75.89$\pm$1.03 67.87$\pm$3.69 60.33$\pm$1.37 84.65$\pm$2.53 86.26$\pm$0.84 49.33$\pm$2.10 83.22$\pm$0.88 52.90$\pm$3.16
O-Emb [46] 74.86$\pm$1.39 75.21$\pm$2.16 44.50$\pm$1.06 86.46$\pm$1.59 77.95$\pm$1.16 65.64$\pm$3.42 79.28$\pm$1.24 64.70$\pm$2.72
@@
Table 1: OOD Detection — Offline Discrimination: AUROC and FPR95 results. Underline and bold denote SOTA among all baselines and all methods, respectively. We report the average results under each setting in the main text, results of each dataset are shown in Table 8 and 9 (Appendix E).
Online Discrimination.
@@
Table 2: OOD Detection — Online Discrimination: Accuracy and Robustness results. We mainly compare our method with embedding-based methods, and bold denotes the best among these methods.
@@
4.3 Scenario II: OOD Quality Estimation
@@
MS-Prob 0.024$\pm$0.020 0.038$\pm$0.020 0.038$\pm$0.018 0.026$\pm$0.018 0.066$\pm$0.015 0.044$\pm$0.016 0.057$\pm$0.018 0.057$\pm$0.022
PPL 0.050$\pm$0.015 0.045$\pm$0.016 0.074$\pm$0.017 0.050$\pm$0.018 0.036$\pm$0.014 0.038$\pm$0.017 0.035$\pm$0.018 0.058$\pm$0.019
I-Emb 0.078$\pm$0.016 0.102$\pm$0.017 0.036$\pm$0.018 0.115$\pm$0.017 0.059$\pm$0.012 0.098$\pm$0.016 0.012$\pm$0.018 0.068$\pm$0.016
O-Emb 0.058$\pm$0.018 0.025$\pm$0.017 0.038$\pm$0.015 0.012$\pm$0.017 0.050$\pm$0.012 0.016$\pm$0.017 0.036$\pm$0.017 0.029$\pm$0.021
@@
Table 3: OOD Quality Estimation: Kendall’s $\tau$ and Spearman correlation between various OOD scores and benchmark quality metric binary matching. Each column shows the correlation when ID and OOD samples are merged. Underline denotes the SOTA among all baselines, and bold denotes the SOTA among our methods. We report the average results under each setting in the main text, results of each dataset are shown in Table 10 and 11 (Appendix E).

### v2

@@ 2.2 Empirical Analysis: Early Stabilization of ID samples
We term this phenomenon “Early Stabilization”: For ID data, GLMs largely complete their reasoning in the mid-to-late stages, and simple adjustments are sufficient after that. However, for OOD data, GLMs can still not complete accurate reasoning at a later stage. They thus can only randomly switch to a specific output pattern, i.e., the scalar mathematical expression pattern. This provides strong evidence that using embedding trajectory for OOD detection may be effective.
3 TV Score: Trajectory Volatility Score for OOD Detection
In Section 2.2, we have identified a significant difference in embedding trajectory volatilities between ID and OOD samples in mathematical reasoning. We now aim to leverage this phenomenon to develop a lightweight OOD detection solution tailored for mathematical reasoning scenarios.
Inspired by static embedding methods that use the ID sample embedding cluster as the reference to measure the Mahalanobis Distance (MaDis) [22], we similarly aim to use the ID sample trajectory cluster as the reference to measure the difference between them and a new sample trajectory. While measuring the difference between an embedding and an embedding cluster only requires a simple MaDis calculation, quantifying the difference between a trajectory and a trajectory cluster is less intuitive. Thus, we seek a transition idea starting from the static space Gaussian assumption.
First, we obey the assumption under static embeddings [3, 43] to fit all ID embeddings at each layer $l$ to a Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\bm{\mu}}_{l},{\bm{\Sigma}}_{l})$. Next, for a new sample with $\bm{y}_{l}$ be its embedding at layer $l$, we map it to its MaDis $f({\bm{y}_{l}}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ with $\mathcal{G}_{l}$ as follows:
$f(\bm{y}_{l})=\left({\bm{y}_{l}}-{\bm{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\bm{y}_{l}}-{\bm{\mu}_{l}}\right)~{}~{}~{}~{}(0\leq l\leq L).$ (4)
Finally, we treat the MaDis differences $\left|f(\bm{y}_{l})-f(\bm{y}_{l-1})\right|$ between adjacent layer-pairs $l-1$ and $l$ as the adjacent-layer volatility of the new sample, and average all adjacent-layer volatilities as the final trajectory volatility score (TV Score $S$) of this sample:
$S=\frac{1}{L}\cdot\sum_{l=1}^{L}\left|f(\bm{y}_{l})-f(\bm{y}_{l-1})\right|.$ (TV Score)
The anticipated trend is that when the two adjacent embeddings after MaDis mapping exhibit a greater difference, the embedding change between the two layers is more volatile compared to ID data. This enables us to identify new samples with trajectory volatility features that significantly deviate from those of the ID samples, thus increasing the likelihood that they are OOD samples.
@@ 3 TV Score: Trajectory Volatility Score for OOD Detection
Furthermore, as a trajectory, outliers in the trajectory may significantly impact feature extraction [21, 28]. To mitigate this, we explore higher-order differential smoothing techniques to enhance trajectory smoothness. We first define the $k$-order embedding $\nabla^{k}\bm{y}_{l}$ and Gaussian distribution $\nabla^{k}\mathcal{G}_{l}=\nabla^{k}\mathcal{N}(\bm{\mu}_{l},\bm{\Sigma}_{l})$ for all $l\leq L-k$ based on the backward difference:
$\nabla^{k}\bm{y}_{l}=\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\bm{y}}_{l+k},~{}~{}~{}\nabla^{k}\mathcal{G}_{l}=\mathcal{N}\left(\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\bm{\mu}}_{l+k},~{}\sum_{i=0}^{k}\mathrm{C}_{k}^{i}{\bm{\Sigma}}_{l+k}\right),$ (5)
where $\mathrm{C}_{k}^{i}=\frac{k!}{i!(k-i)!}$. Similarly, we mapping $\nabla^{k}\bm{y}_{l}$ to its MaDis $\nabla^{k}f({\bm{y}_{l}}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ with $\nabla^{k}\mathcal{G}_{l}$:
$\nabla^{k}f({\bm{y}_{l}})=\left(\nabla^{k}\bm{y}_{l}-\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\bm{\mu}}_{l+k}\right)^{\top}\left(\sum_{i=0}^{k}\mathrm{C}_{k}^{i}{\bm{\Sigma}}_{l+k}\right)^{-1}\left(\nabla^{k}\bm{y}_{l}-\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\bm{\mu}}_{l+k}\right).$ (6)
Following the definition of TV Score, we define the trajectory volatility score after differential smoothing (TV Score w/ DiSmo $\nabla^{k}S$) as follows:
$\nabla^{k}S=\frac{1}{L}\cdot\sum_{l=1}^{L}\left|\nabla^{k}f(\bm{y}_{l})-\nabla^{k}f(\bm{y}_{l-1})\right|.$ (TV Score w/ Dismo)
The Algorithmic Process and Computational Complexity are detailed in Appendix D.
@@ 4.1 Setup
For the ID dataset, we use the MultiArith [44], which consists of Math Word Problems on arithmetic reasoning. For the OOD datasets, we intuitively introduce two types of detection scenarios following [43]: (i) Far-shift OOD scenario, we select the MATH [11] dataset with five domains of algebra, geometry, counting and probability, number theory, and precalculus; (ii) Near-shift OOD scenario, we select five independent datasets: GSM8K [6], SVAMP [39], AddSub [13], SingleEq [18], and SingleOp [18]. We consider the ID data negative(-) and the OOD data positive(+). Refer to Appendix E.1 for basic information and OOD features of these datasets.
@@
Given the limited data size of MultiArith, totaling only 600 samples and lacking a standard division, we allocate 360 samples for training and 240 for testing. However, with such a small test set, randomness in evaluation becomes a concern. To mitigate this, we conduct test sampling and set the sampling size as 1000. Specifically, we denote ID dataset as $\mathcal{D}_{\text{in}}$ and OOD dataset as $\mathcal{D}_{\text{out}}$. For each sampling, the collection is $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$ where $\widetilde{\mathcal{D}}_{\text{out}}\subset\mathcal{D}_{\text{out}}$ and $|\mathcal{D}_{\text{in}}|=|\widetilde{\mathcal{D}}_{\text{out}}|$, this guarantees positive and negative sample balance. We report both the mean and standard variance of the results to enhance the reliability of evaluations. Refer to Appendix E.2 for the ID dataset split.
@@
To measure the application value of our method used in cutting-edge GLMs, we use Llama2-7B [50] and GPT2-XL (1.5B) [5] as our backbones for ID dataset training. Refer to Appendix E.3 for training details. However, there exists uncertainty about the data used in the pre-training phase, especially for Llama2 because its data is closed-source. Some research [57, 59] have confirmed the absence of data leakage in Llama2 for the MATH and GSM8K datasets, we still conduct pre-experiments to examine the rationality of the OOD data selection rigorously. Our criterion is the claim that a dataset can be categorized as OOD if it exceeds the capabilities of the base model, as proposed in prior studies [47, 27]. Results of the pre-experiments are shown in Appendix E.4, and they can confirm that these datasets can be considered as OOD data for the two GLMs.
@@
We compare our method with five training-free baselines where OOD training data are unavailable. We refer to the latest survey [20] to select them for the scarcity of OOD detection methods on GLMs: (1) Maximum Softmax Probability (Prob.) [12]; (2) Monte-Carlo Dropout [8]; (3) Sequence Perplexity [3]; (4) Input Embedding [43]; (5) Output Embedding [43]. Refer to Appendix E.5 for details. Additionally, we set the smoothing order $k$ ranges from 1 to 5 for TV Score w/ DiSmo and report the highest among them when with smoothing.
@@
We divide the OOD detection evaluation into two scenarios: (1) Offline Detection, which classifies whether samples from a given list belong to OOD. For each collection $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$, we report the AUROC [35] and FPR95 metrics. The former represents the area under the ROC curve and the latter represents the value of FPR at 95% TPR. (2) Online Detection, which utilizes the offline detection results to calculate an optimal classification threshold for a direct determination of whether new samples belong to OOD. We introduce the metrics in the corresponding result sub-sections.
4.2 Main Results
Offline Detection.
Table 1 presents the results of the offline detection scenarios.
@@ 4.2 Main Results
In addition, we conduct significant tests (Details are shown in Table 11 - 14). We find that our methods almost pass all significance tests, while the embedding-based methods have the lowest pass rate among baselines, suggesting that their results are more susceptible to sampling error. We also find that the performance of differential smoothing fluctuates greatly in different settings. Therefore, we conduct the ablation of the smoothing order $k$. Results and analyses are shown in Appendix F.1.
Table 1: AUROC and FPR95 results of the Offline Detection scenario. Underline and bold denote SOTA among all baselines and all methods, respectively. We report the average results under each setting in the main text, results of each dataset are shown in Table 11 and 12 (Appendix F).
@@
Max Softmax Prob. [12] 78.66$\pm$1.38 81.44$\pm$3.56 60.14$\pm$1.54 88.91$\pm$2.41 70.54$\pm$1.42 78.29$\pm$2.02 67.12$\pm$1.20 76.27$\pm$2.66
Monte-Carlo Dropout [8] 68.63$\pm$2.21 87.04$\pm$4.88 52.33$\pm$2.21 91.92$\pm$1.89 66.18$\pm$1.87 84.69$\pm$1.65 63.54$\pm$1.72 78.08$\pm$2.50
Perplexity [3] 85.64$\pm$1.46 53.06$\pm$4.36 59.35$\pm$1.89 86.09$\pm$1.89 80.82$\pm$1.04 64.53$\pm$2.10 73.74$\pm$1.12 72.39$\pm$1.27
Input Embedding [43] 75.89$\pm$1.03 67.87$\pm$3.69 60.33$\pm$1.37 84.65$\pm$2.53 86.26$\pm$0.84 49.33$\pm$2.10 83.22$\pm$0.88 52.90$\pm$3.16
Output Embedding [43] 74.86$\pm$1.39 75.21$\pm$2.16 44.50$\pm$1.06 86.46$\pm$1.59 77.95$\pm$1.16 65.64$\pm$3.42 79.28$\pm$1.24 64.70$\pm$2.72
@@
Online Detection.
@@
Table 2: Accuracy and Robustness results of the Online Detection scenario. We mainly compare our method with embedding-based methods, and bold denotes the best among these methods.
@@
5 Generalizability Exploration
5.1 Beyond Detection: OOD Quality Estimation
@@
Table 3: OOD Quality Estimation: Kendall’s $\tau$ and Spearman correlation between various OOD scores and benchmark quality metric binary matching. Each column shows the correlation when ID and OOD samples are merged. Underline denotes the SOTA among all baselines, and bold denotes the SOTA among our methods. We report the average results under each setting in the main text, results of each dataset are shown in Table 13 and 14 (Appendix F).
@@
Max Softmax Prob. 0.024$\pm$0.020 0.038$\pm$0.020 0.038$\pm$0.018 0.026$\pm$0.018 0.066$\pm$0.015 0.044$\pm$0.016 0.057$\pm$0.018 0.057$\pm$0.022
Perplexity 0.050$\pm$0.015 0.045$\pm$0.016 0.074$\pm$0.017 0.050$\pm$0.018 0.036$\pm$0.014 0.038$\pm$0.017 0.035$\pm$0.018 0.058$\pm$0.019
Input Embedding 0.078$\pm$0.016 0.102$\pm$0.017 0.036$\pm$0.018 0.115$\pm$0.017 0.059$\pm$0.012 0.098$\pm$0.016 0.012$\pm$0.018 0.068$\pm$0.016
Output Embedding 0.058$\pm$0.018 0.025$\pm$0.017 0.038$\pm$0.015 0.012$\pm$0.017 0.050$\pm$0.012 0.016$\pm$0.017 0.036$\pm$0.017 0.029$\pm$0.021